Jaimie can bike five times as fast as he can run. To travel 25 km, he needs 2 h more if he is running than if he is biking. What is his average running speed?

Answer by jorel1380(3719)

Let r be Jaimie's running speed. Then his biking speed is 5r. So:
25/r - 2=25/5r
125-10r=25
10r=100
r=10 km/h
Jaimie can run 10 km/h, and bike 50 km/h. ☺☺☺☺